OpenStudy (anonymous):

Write the equation of the line that is parallel to the line 3x − y = −3 and passes through the point (4, −2). y = −one thirdx − 6 y = −one thirdx − 14 y = 3x − 6 y = 3x − 14

OpenStudy (bahrom7893):

parallel to some line means the slopes are the same, so find the slope of the original: 3x-y = -3 -y = -3x-3 y=3x+3

OpenStudy (bahrom7893):

so the slope is 3. Now just use: y-y1 = 3(x-x1) with the point (4, -2) to get the equation of the line.